OpenStudy (anonymous):

In a game of dice, you roll two standard six-sided dice and examine their sum. a) What is the probability of rolling a 7? 1/6 b) If you roll the dice 20 times, what is the probability of rolling a total of 3 sums of 7? c) How many times would you expect to roll the dice before you rolled a sum of 12? 36 I already answered a and c, but I need help with b. If anyone could help, that'd be awesome. Thanks!

OpenStudy (anonymous):

@beccaboo333

OpenStudy (beccaboo333):

Yeah nope. Becca=no good math

OpenStudy (anonymous):

Thanks anyway.

OpenStudy (anonymous):

@Hero can you help?

OpenStudy (anonymous):

@satellite73 can you help?

OpenStudy (anonymous):

Can anyone help me? I have a lot to do and I really need to get through this assignment but I need help.

jimthompson5910 (jim_thompson5910):

have a look at this chart http://bestcase.files.wordpress.com/2011/01/dicediagram.jpg it shows all 36 ways to roll 2 dice

jimthompson5910 (jim_thompson5910):

the cells represent the sums of the two dice (from the row & and column headers)

jimthompson5910 (jim_thompson5910):

oh wait, you already have the answer of 1/6, nvm but you would could the number of 7s that show up (6 of them) and divide by 36 to get 6/36 = 1/6 So part a) is correct

OpenStudy (anonymous):

Alright what about b?

jimthompson5910 (jim_thompson5910):

to answer part b, we need to use a binomial distribution

jimthompson5910 (jim_thompson5910):

in this case, probability of success (ie rolling a 7) is p = 1/6 n = 20 (ie we have 20 trials or rolls) k = 3 (we want exactly 3 rolls of 7) and we plug all that into (n C k)*(p)^(k)*(1-p)^(n-k)

jimthompson5910 (jim_thompson5910):

the notation n C k means "n choose k" and can be found using the formula n C k = (n!)/(k!*(n-k)!)

jimthompson5910 (jim_thompson5910):

let me know if that helps at all

OpenStudy (anonymous):

Sorry, I had to go eat. @jim_thompson5910 so it would be: n C k = (20)/3 * (20 - 3)) = 0.392 (0.392) * (1/6)^(3) * (1 - (1/6))^(20 - 3) = 1.8155E - 03 = 0.0018155 Is that right?

jimthompson5910 (jim_thompson5910):

20! and not just 20

jimthompson5910 (jim_thompson5910):

the exclamation mark is factorial notation (sorry I should have stated that)

jimthompson5910 (jim_thompson5910):

so 20! = 20*19*18...*3*2*1 most calculators have a factorial button so you don't have to type in all that

jimthompson5910 (jim_thompson5910):

the same is true for the 3 and the (20-3)

OpenStudy (anonymous):

finally someone calls it dice i am sick to death of "number cubes"

OpenStudy (anonymous):

Ohhh... I don't know how to do that on a calculator o.o

OpenStudy (anonymous):

do it in one step in wolfram

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have?

OpenStudy (anonymous):

Casio fx-9750GII but I have no clue how to like do anything on it. I know how to do very little on it.

jimthompson5910 (jim_thompson5910):

ok let me bring up a users manual on it the images I'm looking at suggest that this thing is similar to a ti-83 (which I am familiar with)

jimthompson5910 (jim_thompson5910):

ok found it

OpenStudy (anonymous):

Okay thank you :) Yeah, it's kinda like a ti but different.

jimthompson5910 (jim_thompson5910):

you hit the option button, then you find the PROB submenu

jimthompson5910 (jim_thompson5910):

in there you'll see n P r (for permutations) and n C r (for combinations) we'll use n C r

OpenStudy (anonymous):

Okay found it

jimthompson5910 (jim_thompson5910):

most likely (based on how the ti83 works), you type it in like this 20 n C r 3

jimthompson5910 (jim_thompson5910):

so you type 20 first, bring up the n C r, then you type 3 to finish it off

OpenStudy (anonymous):

Alright, I did that and the number 1140 came up

jimthompson5910 (jim_thompson5910):

good, that means 20 C 3 = 1140

jimthompson5910 (jim_thompson5910):

So (n C k)*(p)^(k)*(1-p)^(n-k) (20 C 3)*(1/6)^(3)*(1-1/6)^(20-3) (1140)*(1/6)^(3)*(1-1/6)^(20-3) and I'll let you compute the rest

OpenStudy (anonymous):

I got 5.277777778 Is that right?

jimthompson5910 (jim_thompson5910):

that's way too big, remember probabilities are between 0 and 1

jimthompson5910 (jim_thompson5910):

0 meaning 0% (event never happens) 1 meaning 100% (event is guaranteed to happen)

OpenStudy (anonymous):

Oh, doesn't that mean I need to divide it by like 100 or something?

jimthompson5910 (jim_thompson5910):

not quite, it means that you know what window your answer should be in but the answer isn't 5%

jimthompson5910 (jim_thompson5910):

what are you getting when you compute (1/6)^(3)

OpenStudy (anonymous):

4.6296E - 03

jimthompson5910 (jim_thompson5910):

so that is also written as 0.0046296

jimthompson5910 (jim_thompson5910):

how about (1-1/6)^(20-3)

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

just 1?

jimthompson5910 (jim_thompson5910):

it should be some decimal number

jimthompson5910 (jim_thompson5910):

so this is probably where the error is happening

OpenStudy (anonymous):

probably, because it just came up as 1

jimthompson5910 (jim_thompson5910):

what did you type in exactly?

jimthompson5910 (jim_thompson5910):

you should have typed in (1-1/6)^(20-3) exactly as you see it

jimthompson5910 (jim_thompson5910):

oh and I should mention that 1-1/6 is not 0 it's really 1 - (1/6) so you could simplify that portion to get 1-1/6 = 5/6

OpenStudy (anonymous):

oh I think I did something wrong before because I just did it again and this time I got 0.04507

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so (1140)*(1/6)^(3)*(1-1/6)^(20-3) turns into (1140)*(0.0046296)*(0.04507)

jimthompson5910 (jim_thompson5910):

of course, you can type in all of (1140)*(1/6)^(3)*(1-1/6)^(20-3) to get the direct answer (but it helps to break it down a bit)

OpenStudy (anonymous):

I got 0.23787 Does that look right?

jimthompson5910 (jim_thompson5910):

when I typed in (1140)*(1/6)^(3)*(1-1/6)^(20-3), I'm getting 0.23788656613786 So it does look right

jimthompson5910 (jim_thompson5910):

So you roughly have a 23.7% chance of rolling exactly 3 sevens

jimthompson5910 (jim_thompson5910):

or 23.8% I mean (if you're rounding percentages to 1 decimal place)

OpenStudy (anonymous):

Omg yay! Thank you so much! I have a couple of other problems on this assignment that I could use some help with, if you don't mind could you help? If not, that's okay. But seriously, thank you.

jimthompson5910 (jim_thompson5910):

sure, go ahead

OpenStudy (anonymous):

Should I put them in a new question or just put them on here? And thank you!

jimthompson5910 (jim_thompson5910):

either way works, a new question is probably best so things don't get too cluttered or laggy

OpenStudy (anonymous):

Alright. One sec.

jimthompson5910 (jim_thompson5910):

alright